OpenStudy (anonymous):

rationalize the denominator 14sqrt3/3sqrt7

OpenStudy (ashish56):

\[14\sqrt{3}/3\sqrt{7}=14\sqrt{3}\times \sqrt{7}/3\times(\sqrt{7})^{2}=14\sqrt{21}/21\] =\[2\sqrt{21}/3\]